# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v3

@@ VIII Discussion and perspectives
The quest for a time operator is a fundamental question in Quantum Mechanics Aharonov and Bohm (1961); Susskind and Glogower (1964); Unruh and Wald (1989), which is still open Höhn et al. (2021). In our case, the time operator emerges from the quantum fluctuations of the global time origin, in close analogy with the fluctuations of the global phase of a condensate. Indeed, in the $(t,\phi)$ formalism, the time operator emerges from the fluctuations of the global phase shift $\phi_{0}$ in the periodic motion of the spontaneous Floquet state. The connection between time and phase has been pointed out since early approaches to the problem Susskind and Glogower (1964) because any realization of a time operator is expected to be based on the oscillation of a clock. The time-phase analogy suggests that spontaneous Floquet states should be then macroscopically excited states, with an energy well above the ground state. As a result, the CES state provides a condensed-matter setup where a tangible realization of a time operator can be studied.
@@
Our work also finds a major application in the field of time crystals, since a spontaneous Floquet state represents a specific realization of a continuous time crystal de Nova and Sols (2022). Indeed, some results of our work are connected with the time-crystal literature. For instance, the role of the generalized Gibbs ensemble in dissipative time crystals was thoroughly discussed in Ref. Booker et al. (2020). The emergence of continuous time crystals from excited eigenstates, in the same fashion as spontaneous Floquet states, was predicted in Ref. Syrwid et al. (2017). Time supersolids, like the CES state, were originally observed in Ref. Autti et al. (2018), to which the formalism of this work is directly applicable. On the other hand, Floquet thermodynamics can provide a novel tool for the characterization of both continuous and discrete time crystals. Moreover, the analysis presented in Sec. IV suggests that certain periodic systems identified as time crystals may not genuinely break continuous time-translation symmetry since they can arise from constant motion along a closed orbit generated by the broken-symmetry transformations. Thus, the presence of a genuine temporal FNG mode should be regarded as the characteristic signature of any bona fide continuous time crystal.
@@
Although we have used atomic condensates to illustrate our Floquet formalism, it can be straightforwardly extrapolated to magnonic condensates, nonlinear optical fibers or quantum fluids of light. In general, spontaneous Floquet states extend the field of nonlinear Floquet waves Kreil et al. (2019); Träger et al. (2021) to scenarios without external driving. Spontaneous Floquet states in fermionic systems described by HF approaches, such as quantum Hall states, are also plausible de Nova and Sols (2022). In addition, self-consistent BCS-type theories are known to display periodic oscillations in momentum space Barankov et al. (2004); Foster et al. (2014); Perfetto and Stefanucci (2020).
@@
Future extensions of the present work should include finite temperature effects and quantum corrections beyond linearity, both for stationary and Floquet states exhibiting simultaneous symmetry breaking. Of particular interest is the general study of the nonperturbative quantum diffusion of the spontaneously broken symmetries, generalizing the works of Refs. Lewenstein and You (1996); Dziarmaga (2004). In the case of the temporal Floquet-Nambu-Goldstone mode, this could lead to suggestive concepts such as time diffusion or time eigenstates, in analogy with the phase of a condensate Leggett and Sols (1991); Sols (1994). A subject of ongoing research de Nova et al. is the achievement of a fully quantum time crystal from the CES state that spontaneously breaks time-translation symmetry, without phase locking its time origin as in the present work. This can be done by either approaching the GS-CES phase transition, where quantum fluctuations become critical, or by starting from a quantum black-hole laser de Nova and Sols (2023).
Acknowledgements.
@@ A.1 Elliptic functions
$\textrm{sn}(u,\nu)\equiv\sin\left(\textrm{am}(u,\nu)\right),~{}\textrm{cn}(u,\nu)\equiv\cos\left(\textrm{am}(u,\nu)\right),$ (175)
@@
$F(\phi,\nu)\equiv\int^{\phi}_{0}\frac{\mathrm{d}\varphi}{\sqrt{1-\nu\sin^{2}\varphi}},~{}K(\nu)\equiv F\left(\frac{\pi}{2},\nu\right).$ (177)
As a consequence of these relations, $\text{sn}(u+2K(\nu),\nu)=-\text{sn}(u,\nu),~{}\text{cn}(u+2K(\nu),\nu)=-\text{cn}(u,\nu)$, and thus $\text{sn}(u,\nu),\,\text{cn}(u,\nu)$ are periodic functions with period $4K(\nu)$. The incomplete and complete elliptic integrals of the second kind are
$E(\phi,\nu)\equiv\int_{0}^{\phi}\mathrm{d}\varphi~{}\sqrt{1-\nu\sin^{2}\varphi},~{}E(\nu)\equiv E\left(\frac{\pi}{2},\nu\right),$ (178)
@@
$G_{2n}(\nu)$$\equiv$ $\int_{0}^{\frac{\pi}{2}}\sin^{2n}\phi\sqrt{1-\nu\sin^{2}\phi}~{}\mathrm{d}\phi,$
which appear when evaluating the particle number or the energy of a cnoidal wave. They can be shown to obey the inter-recursive relations
@@
with $F_{0}(\nu)=K(\nu),~{}G_{0}(\nu)=E(\nu)$.
@@ A.2 GPG solutions in a ring
$n^{\prime 2}=4(n-n_{1})(n-n_{2})(n-n_{3}),~{}\theta^{\prime}(x)=v(x)=\frac{J}{n(x)},$ (186)
where the three densities $n_{i}$, $i=1,2,3$, are the roots of the polynomial
@@
After integrating the equation for $n(x)$ in Eq. (186) and then that of $\theta(x)$, we arrive at the cnoidal wave of Eq. (85). From its expression, we can compute the pure cnoidal contribution to the momentum
@@
Moreover, $|\Psi_{0}(x)|^{2}=|\Psi_{\mathbf{n}}(x)|^{2}$, and thus the total particle number is directly obtained from the integral of the cnoidal density. With the help of these results, one arrives at Eq. (93). The trick to solve that system of equations is to express all roots $n_{i}$ in terms of $\nu$, $n_{i}=n_{i}(\nu)$. For that purpose, one combines the first two lines of Eq. (93) to express $n_{1},n_{3}$ in terms of $\nu$, and invokes the definition of $\nu$ itself, Eq. (86), to obtain $n_{2}(\nu)$. Inserting these results into the third line of Eq. (93) yields a single equation for $\nu$, which can be easily numerically solved as $0\leq\nu\leq 1$. Once $\nu$ is obtained, the densities $n_{i}=n_{i}(\nu)$ are computed, and the complete GPG wavefunction and the total energy are determined.
@@
$F_{AB}=\int\mathrm{d}\mathbf{x}~{}[\partial_{A}\eta\,\partial_{B}n-\partial_{A}n\,\partial_{B}\eta].$ (191)

### v4

@@ 8 Discussion and perspectives
The quest for a time operator is a fundamental problem in Quantum Mechanics [131, 132, 133, 134], which is still open [135]. In our case, the time operator emerges from the quantum fluctuations of the global time origin, in close analogy with the fluctuations of the global phase of a condensate. Indeed, in the $(t,\phi)$ formalism, the time operator emerges from the fluctuations of the global phase shift $\phi_{0}$ in the periodic motion of the spontaneous Floquet state. The connection between time and phase has been pointed out since early approaches to the subject [132] because any realization of a time operator is expected to be based on the oscillation of a clock. The time-phase analogy suggests that spontaneous Floquet states should be then macroscopically excited states, with an energy well above the ground state. As a result, the CES state provides a condensed-matter setup where a tangible realization of a time operator can be studied. In general, each specific proposal of time operator presents its own characteristic features, determined by the underlying system acting as a clock, such as those based on kaons [136, 137, 138] or the so-called time of arrival [139, 140, 141, 142].
@@
Our work also finds a major application in the field of time crystals, since a spontaneous Floquet state represents a specific realization of continuous time crystal [43]. Indeed, some results of our work are connected with the time-crystal literature. For instance, the role of the generalized Gibbs ensemble in dissipative time crystals was thoroughly discussed in Ref. [143]. The emergence of continuous time crystals from excited eigenstates, in a similar fashion to spontaneous Floquet states, was predicted in Ref. [144]. Time supersolids, like the CES state, were originally observed in Ref. [27], to which the formalism of this work is directly applicable. Our explicit characterization of the temporal FNG mode complements other field-theoretical examples present in the literature [145, 146, 147, 148]. On the other hand, Floquet thermodynamics can provide a novel tool for the characterization of both continuous and discrete time crystals. Moreover, the analysis presented in Sec. 4 suggests that certain periodic systems identified as time crystals may not genuinely break continuous time-translation symmetry since they can arise from constant motion along a closed orbit generated by other broken-symmetry transformations. As a result, there is not a truly temporal NG mode since this can be expressed in terms of the remaining NG modes. Thus, the presence of a genuine temporal FNG mode should be regarded as the characteristic signature of any bona fide continuous time crystal.
@@
Extended Gross-Pitaevskii equations are also expected to exhibit spontaneous Floquet states. Although we have used atomic condensates to illustrate our Floquet formalism, this can be straightforwardly extrapolated to magnonic condensates, nonlinear optical fibers or quantum fluids of light. Spontaneous Floquet states in fermionic systems described by Hartree-Fock approaches, such as quantum Hall states, are also plausible [43]. In addition, self-consistent BCS-type theories are known to display periodic oscillations in momentum space [149, 150, 151]. Spontaneous Floquet states can also arise in few-body [152] and many-body [29] dissipative quantum systems whose effective dynamics is described by self-consistent equations of motion. In general, spontaneous Floquet states extend the field of nonlinear Floquet waves [153, 154] to scenarios without external driving.
@@
Future extensions of the present work should include finite-temperature effects and quantum corrections beyond linearity, both for stationary and Floquet states exhibiting simultaneous symmetry breaking. Of particular interest is the general study of the nonperturbative quantum diffusion of the spontaneously broken symmetries, generalizing the works of Refs. [55, 56]. In the case of the temporal Floquet-Nambu-Goldstone mode, this could lead to suggestive concepts such as time diffusion or time eigenstates, in analogy with the phase [155, 156]. A fully quantum time crystal based on a spontaneous Floquet state, spontaneously breaking time-translation symmetry without phase-locking its time origin as in the present work, can be achieved through the self-amplification of quantum Hawking radiation [157].
@@ A.1 Elliptic functions
$\textrm{sn}(u,\nu)$ $\equiv$ $\sin\left(\textrm{am}(u,\nu)\right),$ (184)
$\textrm{cn}(u,\nu)$ $\equiv$ $\cos\left(\textrm{am}(u,\nu)\right),$
@@
$F(\phi,\nu)$ $\equiv$ $\int^{\phi}_{0}\frac{\mathrm{d}\varphi}{\sqrt{1-\nu\sin^{2}\varphi}},$
$K(\nu)$ $\equiv$ $F\left(\frac{\pi}{2},\nu\right).$ (186)
As a consequence of these relations, $\text{sn}(u+2K(\nu),\nu)=-\text{sn}(u,\nu)$, $\text{cn}(u+2K(\nu),\nu)=-\text{cn}(u,\nu)$, and thus $\text{sn}(u,\nu),\,\text{cn}(u,\nu)$ are periodic functions with period $4K(\nu)$.
The incomplete and complete elliptic integrals of the second kind are
$E(\phi,\nu)$ $\equiv$ $\int_{0}^{\phi}\mathrm{d}\varphi\leavevmode\nobreak\ \sqrt{1-\nu\sin^{2}\varphi}\,$, (187)
$E(\nu)$ $\equiv$ $E\left(\frac{\pi}{2},\nu\right),$
@@
$G_{2n}(\nu)$ $\equiv$ $\int_{0}^{\frac{\pi}{2}}\sin^{2n}\phi\sqrt{1-\nu\sin^{2}\phi}\leavevmode\nobreak\ \mathrm{d}\phi,$
which appear when evaluating the particle number and energy of a cnoidal wave. They can be shown to obey the inter-recursive relations
@@
with $F_{0}(\nu)=K(\nu),\leavevmode\nobreak\ G_{0}(\nu)=E(\nu)$.
Interestingly, by using elliptic functions, it is straightforward to solve analytically the textbook problem of the motion of a pendulum of mass $m$ subject to a constant gravitational field $g$:
$mL\ddot{\theta}=-mg\sin\theta,$ (192)
$\theta$ being the angle of separation with respect to the equilibrium position. Energy conservation reads
$\frac{1}{2}mL^{2}\dot{\theta}^{2}+mgL(1-\cos\theta)=E,$ (193)
which can be rewritten as
$\frac{1}{2}\dot{\theta}^{2}+2\frac{g}{L}\sin^{2}\frac{\theta}{2}=e\equiv\frac{E}{mL^{2}}.$ (194)
Integrating this equation yields
$\theta(t)$ $=2\arcsin\left(\sqrt{\nu}\leavevmode\nobreak\ \textrm{sn}\left[\sqrt{\frac{g}{L}}(t-t_{0}),\nu\right]\right),\leavevmode\nobreak\ \nu\leq 1,$
$\theta(t)$ $=2\,\textrm{am}\left[\sqrt{\frac{e}{2}}(t-t_{0}),\frac{1}{\nu}\right],\leavevmode\nobreak\ \nu>1,$ (195)
$\nu$ $\equiv\frac{eL}{2g},$
with $t_{0}$ a time-shift determined by the initial condition. The first line describes the bounded oscillation of a pendulum while the second line represents its full circular motion with an energy larger than that of the unstable equilibrium position ($\theta=\pi$). The corresponding periods of motion are
$T$ $=$ $4\sqrt{\frac{L}{g}}K(\nu),\leavevmode\nobreak\ \nu\leq 1,$
$T$ $=$ $\sqrt{\frac{8}{e}}K\left(\frac{1}{\nu}\right),\leavevmode\nobreak\ \nu>1.$ (196)
The usual limit of small sinusoidal oscillation is recovered for $\nu\to 0$, where
$\theta(t)\simeq\sqrt{\frac{2eL}{g}}\sin\omega_{0}(t-t_{0}),\leavevmode\nobreak\ \omega_{0}=\sqrt{\frac{g}{L}}.$ (197)
@@ A.2 GPG solutions in a ring
$n^{\prime 2}$ $=$ $4(n-n_{1})(n-n_{2})(n-n_{3}),$
$\theta^{\prime}(x)$ $=$ $v(x)=\frac{J}{n(x)},$ (201)
where the three densities $0\leq n_{1}\leq n_{2}\leq n_{3}$ are the roots of the polynomial
@@
After integrating the equation for $n(x)$ in the first line of Eq. (A.2) and then that for $\theta(x)$ in the second line, we arrive at the cnoidal wave of Eq. (89). From its expression, we can compute the pure cnoidal contribution to the momentum
@@
Moreover, $|\Psi_{0}(x)|^{2}=|\Psi_{\mathbf{n}}(x)|^{2}$, and thus the total particle number is directly obtained from the integral of the cnoidal density. With the help of these results, one arrives at Eq. (97). The trick to solve that system of equations is to express all roots $n_{i}$ in terms of $\nu$, $n_{i}=n_{i}(\nu)$. For that purpose, one combines the first two lines of Eq. (97) to express $n_{1},n_{3}$ in terms of $\nu$, and invokes the definition of $\nu$ itself, Eq. (90), to obtain $n_{2}(\nu)$. Inserting these results into the third line of Eq. (97) yields a single equation for $\nu$, which can be easily numerically solved as $0\leq\nu\leq 1$. Once $\nu$ is obtained, the densities $n_{i}=n_{i}(\nu)$ are evaluated, and the complete GPG wavefunction and the total energy are determined.
@@
$F_{AB}=\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ [\partial_{A}\eta\,\partial_{B}n-\partial_{A}n\,\partial_{B}\eta].$ (206)